Aslam's age is half of his father's age, but 15 years ago his age was just 1/3 of his father's age. Find out his present age
Answer by TimothyLamb(4379)

x = aslam age
y = father age
---
y = 2x
y - 15 = 3(x - 15)
---
put the system of linear equations into standard form
---
2x - y = 0
y - 15 = 3x - 45
---
2x - y = 0
3x - y = 30
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = aslam age = 30
y = father age = 60
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---